### v2

# Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning

Abstract
Prompt tuning methods have shown comparable performance to general training methods as parameter-efficient fine-tuning (PEFT) methods in various natural language understanding tasks. However, existing prompt tuning methods still utilize the entire model architecture even when solving a specific task, which prevents them from accelerating inference speed during the application procedure. In this paper, we propose a novel prompt tuning framework called Skeleton to efficiently utilize a language model in terms of memory and time complexity for solving various tasks, retaining only task-relevant neurons by using an explainability method. From our framework, we can efficiently solve various tasks by using only task-relevant neurons and prepending adequate task-specific prompt tokens with only a single language model. Experiments reveal that our method significantly enhances inference efficiency (at most ×$\times 1.73$ speed up) for various widely used benchmarks, showing comparable performance to the prompt tuning method. Moreover, our method is applicable across various transformer-based architectures, confirming its practicality and scalability.
Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning
Nakyeong Yang1, Jiwon Moon1, Junseok Kim2, Yunah Jang1 and Kyomin Jung1 1Seoul National University, 2Pohang University of Science and Technology {yny0506, wldnjs913, vn2209, kjung}@snu.ac.kr {junseokkim00}@postech.ac.kr
1 Introduction
With the success of parameter-efficient fine-tuning (PEFT) methods, pre-trained language models perform various tasks by adjusting only small numbers of parameters. Among PEFT methods, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) are widely used to solve multiple tasks since they require only a few tunable tokens to be prepended to the input tokens while freezing the original parameters of a language model (Chan et al., 2023; Huang et al., 2024). This strategy is memory efficient, as it enables a model to solve various tasks with only a few task-specific prompt tokens while retaining the original parameters. However, we can raise one research question regarding this task-specific trait of prompt tuning:
"Leveraging prompt tuning, could we condense task knowledge into a few task neurons in a frozen model to efficiently utilize it?"
Figure 1: Skeleton Framework Overview. Our framework activates only task-localized paths of language models in the prompt tuning procedure. Each color (i.g., red, blue, and yellow) corresponds to a specific task; we obtain a task-specific prompt and neuron indices after executing our framework. Gray-colored paths in the figure represent task-irrelevant paths that are never used during training and inference, thereby improving inference speed.
Existing prompt tuning frameworks are limited by the need to use the full original model architecture during inference. In fact, these methods may require additional inference time, as the prepended prompt tokens introduce further computational overhead. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) have attempted to accelerate the inference speed of prompt-tuned models by pruning unnecessary prompt tokens. However, they have only marginally improved inference efficiency, as the number of pruned tokens is a small fraction ($\leq 0.1\%$) of all the parameters of a language model (Lester et al., 2021).
Prior studies have attempted to investigate and analyze task-relevant neurons in language models to explain the language models’ behavior (Panigrahi et al., 2023; Yang et al., 2024). Specifically, Panigrahi et al. (2023) has suggested training new parameters, which mask the original parameters of language models to determine task neurons. Yang et al. (2024) has utilized an explainability method to determine task-relevant neurons and reveal that some task-relevant neurons significantly affect the outputs of specific tasks (i.e., natural language understanding, social bias). However, while they have introduced a method to identify task-relevant neurons, they fall short of proposing a practical approach to efficiently utilize those neurons during training and inference.
To address this challenge, we propose a novel training framework, Skeleton, designed to enhance the efficiency of language models by leveraging only task-relevant neurons for solving various tasks. Specifically, we use attribution (Yang et al., 2023, 2024), an explainability method, to identify and utilize only task-relevant neurons in the prompt tuning procedure. Our framework requires only a few prompt tokens to be updated and freezes the original parameters of a language model, as it adopts the prompt tuning method. Furthermore, Skeleton dramatically accelerates the inference speed of language models since it utilizes only task-relevant neurons. For example, in order to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only the task-relevant neurons to solve the task.
We demonstrate our method in various widely used natural language understanding benchmarks and reveal that our method significantly outperforms existing PEFT methods in inference efficiency without sacrificing the original performance of each task. Surprisingly, our method improves the inference speed up to ×$\times 1.73$ with 65% of the parameters pruning rate. We additionally conduct various analyses to offer a guideline for using our new prompt tuning framework, such as investigating which types of layers are crucial for processing task knowledge in a language model. Furthermore, our prompt tuning framework is practical and scalable since it is applicable to any transformer-based architecture. In summary, this work makes the following contributions:
  • We first define task neuron and propose Task Neuron Localized Prompt Tuning (Skeleton), a practical training framework for accelerating the inference speed of a language model for various tasks.
  • We reveal that Skeleton successfully enhances the inference speed of each task without compromising the task performance.
2 Backgrounds
2.1 Parameter-efficient Fine-tuning
Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023). Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various NLP tasks (Wang et al., 2023; Bansal et al., 2023). Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model. However, they still face a limitation as they are not parameter-efficient during the inference process. Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) propose prompt token selection methods to enhance the inference speed of prompt-tuned language models. Nevertheless, they have only marginally enhanced inference efficiency since the number of pruned tokens is only a tiny proportion of the entire number of input tokens.
2.2 Task Knowledge Localization
In spite of the significant performance of language models, it remains challenging to accurately investigate the role of each parameter in language models in dealing with a specific task. Therefore, existing studies have attempted to identify task-specific neurons in executing a specific task (Panigrahi et al., 2023; Yang et al., 2023, 2024). Panigrahi et al. (2023); Yang et al. (2023, 2024) have studied to localize neurons for performing a specific task by clarifying the role of each parameter in a language model. Panigrahi et al. (2023) has suggested a training-based task knowledge localization method, model grafting, which localizes a task knowledge by training new parameters to mask original parameters. Although it has effectively identified task neurons, it has required excessive additional storage for newly trained masking parameters, which is equivalent to the number of model parameters. Yang et al. (2023, 2024) has illuminated task neurons by utilizing the attribution technique (Shrikumar et al., 2016), an explainability method that derives the importance of each feature when solving a specific task. Yang et al. (2023, 2024) has verified that the attribution effectively detects task neurons for solving a specific task and proposed a task neuron detection method applicable to language modeling tasks. However, Yang et al. (2023, 2024) has a limitation in just pruning the task neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded. Furthermore, the prior study has only proposed the method for pruning task neurons and is far from proposing a practical method to improve the inference speed of task-localized language models. The architecture of transformer variants consists of complex neural computation processes; thus, satisfying the integrity of dimension dependency for adjacent networks while pruning the weights is challenging. We adopt the attribution method (Yang et al., 2023, 2024) and propose a new practical framework for accelerating a language model, where the framework utilizes only task neurons while training a language model via prompt tuning, resolving the challenges in pruning networks in transformer variants.
3 Problem Definition
3.1 Prompt Tuning
We adopt the prompt tuning method to propose a new framework for accelerating a language model. Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)=\prod_{k=1}^{K}\mathcal{P}_{\theta}(y_{k}|x,y_{1},...,y_{k-1})$ that represents a language model, where $\theta$ is parameters of the language model. Prompting is the method of adding extra information to the model to condition during its generation of $y$. Normally, prompting is achieved by adding a series of tokens $P_{\tau}\in\mathbb{R}^{l\times d}$ to the input $x$ as $\mathcal{P}_{\theta}(y|[P_{\tau};x])$, where $l$ and $d$ are token length and feature dimension, and $P_{\tau}$ is trainable parameters for a specific task $\tau$. In the prompt tuning process, we update only $P_{\tau}$ and get $P^{\prime}_{\tau}$ to maximize the likelihood of $y$ using gradient descent while freezing the parameters of the language model $\theta$. After the prompt tuning, we get the new conditional generation function $\mathcal{P}_{\theta}(y|[P^{\prime}_{\tau};x])$.
3.2 Skeleton Framework
In the Skeleton framework, we still update only prompt tokens, freezing the original parameters. However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we first define task neuron as follows:
Task Neuron.
Let $\mathcal{P}_{\theta}$ be a language model, where $\theta=\{\theta_{1},...,\theta_{N}\}$ is a set of parameters of it. Then, $\theta_{\tau}\subset\theta$ is defined as task neurons for a task $\tau$ if the below equation is satisfied:
$\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta}([P^{\prime}_{\tau};x]))\approx\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta_{\tau}}([P^{\prime\prime}_{\tau};x]))$ (1)
where $\mathcal{D}_{\tau}$ means a dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy). $P^{\prime}_{\tau}$ and $P^{\prime\prime}_{\tau}$ are task-specific prompts. $N$ means the total number of parameters in a language model.
Skeleton framework adopts the prompt tuning strategy using only task neurons $\theta_{\tau}$, and get the new task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime\prime}_{\tau};x])$. Therefore, we can utilize the language model more time-efficiently by using only those neurons within the Skeleton framework. In addition to being time efficient, this framework is also memory-efficient since it requires only a small fraction of original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime\prime}_{\tau}$.
4 Methods
In this section, we describe the process of Skeleton. We first train a language model using the prompt tuning method to inject prior knowledge of a specific task into it. After that, we quantify the task relevance of each neuron to determine task neurons in a language model. We finally retrain the prompt tokens with the prompt tuning method to condense the task knowledge to only the task neurons. Note that the original language model maintains only task neurons via a pruning method in this process. After applying Skeleton framework, we can utilize the language model efficiently by using only the task-specific prompt and task neurons, accelerating the inference speed for a specific task. The entire process of Skeleton is described in Figure 2 and Algorithm 1.
Figure 2: Task Neuron Localization Process of Skeleton. After training a model using our framework, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure. Red-colored values in the prune weights, $\tilde{W}_{1}$ and $\tilde{W}_{2}$, mean task neurons for a specific task (task A). The weight of FFN2 is pruned depending on FFN1’s dimension, and the output representation of FFN2 is zero-recovered (Blue-colored values) to match the output dimension of the residual network (described in Section 4.5).
4.1 Quantifying Task Relevance of Neurons
This section describes the process of computing the task relevance of each neuron. We adopt an attribution method (Yang et al., 2023, 2024) to quantify the skill relevance of neurons in the pre-trained language models. Formally, given a task-specific prompt tokens $P_{\tau}$ and text input tokens $x$ provided to $\mathcal{P}_{\theta}$, the contribution of the $i$-th neuron to the prediction of an output $y$ is defined as follows:
$A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$ (2)
where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])/\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron.
We adopt transformer variants; thus, activations and gradients are computed for all input token representations. Therefore, if an input text $x_{j}$ includes a $K_{j}$ number of tokens, each neuron has a $K_{j}$ number of attributions. In addition, there are multiple data instances for each task $\tau$; thus, we aggregate attributions for tokens and instances as follows:
$A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)=\sum_{j}^{N}\sum_{k}^{K_{j}}\frac{1}{K_{j}}A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ (3)
where $A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ means the attribution scores computed for an input text token $t_{k}\in x_{j}$. $\mathcal{D}^{\tau}$ and $N$ mean a task-specific dataset and the number of instances in the dataset, respectively. Therefore, $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ is the attribution score computed for a specific task dataset $\mathcal{D}^{\tau}$. Although $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ can be computed using the entire dataset, we report the experimental results of computing it using only a significantly small amount of data (e.g., only twenty data samples) to ensure the efficiency of our method. To simplify the notation, $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ will be abbreviated as $A_{i}$ from this point forward.
4.2 Identifying Task-relevant Neurons.
After quantifying each neuron’s task relevance, we determine task-relevant neurons and eliminate task-irrelevant neurons using a structured pruning method. We first sort the neurons of the whole target layers by the task relevance scores. Then, we determine the optimal neuron ratio using the Binary search algorithm to retain as few task neurons as possible without compromising the task knowledge. Specifically, we compute the accuracy for the validation set $\mathcal{D}^{val}$ to search for the optimal neuron ratio that does not degrade the task performance. We define a margin $\psi$ and regard $\theta_{\tau}$ as task neurons if the accuracy degradation after neuron localization is smaller than the specified margin $\psi$.
4.3 Task Neuron Localization.
After identifying task neurons, we eliminate task-irrelevant neurons from a language model using a structured pruning method. Suppose that a weight matrix $W\in\mathbb{R}^{d\times m}$ is a linear matrix multiplication parameter, and then the matrix after pruning is denoted as $\tilde{W}=(W_{ij})_{\begin{subarray}{c}1\leq i\leq d\\ j\in\mathcal{M}\end{subarray}}$, where $\mathcal{M}$ is the set of task-relevant neuron indices about the $W$. If the bias term $b\in\mathbb{R}^{m}$ is added to the operation for an affine transformation, the bias term can also be pruned by performing the $\tilde{b}=(b_{i})_{i\in\mathcal{M}}$ operation similarly. The task-localized parameters are used to compute the new representation by performing the transformation operation $h\tilde{W}$ or $h\tilde{W}+\tilde{b}$. Notice that this method is model-agnostic since all neural network models consist of linear transformation layers. For example, transformer variants have self-attention and feed-forward network (FFN) modules, all of which include linear matrix multiplication operations. After localizing a language model, we further intensify task knowledge of a localized language model using a prompt tuning method.
4.4 Task Neuron Localized Prompt Tuning
Our Skeleton framework follows the four steps to train a model: (1) the first training step for injecting prior task knowledge; for injecting prior task knowledge; (2) the quantification step of computing task relevance for each neuron; (3) the task neuron localization step; (4) the final prompt tuning step for condensing task knowledge.
The Prior Knowledge Injection Step.
The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a specific task in order to accurately quantify the task-specific knowledge of each neuron. Therefore, we first execute the prior knowledge injection process. Specifically, we train a language model $\mathcal{P}_{\theta}$ via prompt tuning and get a task-specific prompt $P_{\tau}^{\prime}$.
The Quantification Step.
After injecting task knowledge into the language model, we quantify the task relevance $A_{i}$ of each neuron for $\mathcal{P}_{\theta}$ and $P_{\tau}^{\prime}$ using Equation 2 and 3.
The Task Neuron Localization Step.
After quantifying the task relevance $A_{i}$, we first sort the neurons by the task relevance in descending order. Then, we select and maintain only the top-$p$ neurons $\theta_{\tau}\subset\theta$ from the language model by pruning task-irrelevant neurons (Section 4.3). Note that we use a Binary search algorithm to find the optimal ratio of neurons $p\in[0.0,0.05,...,0.95,1.0]$. This search process requires only a maximum of five trials since we search pruning ratio in 0.05 units.
The Knowledge Condensation Step.
During the task neuron localization step, we condense the task knowledge into only the task neurons. Although we never update the original parameters of the language model, we can condense the task knowledge into the task neurons by updating only $P_{\tau}^{\prime}$. We train the task neuron localized model $P_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$. The Binary search algorithm determines the optimal pruning rate after this step.
Algorithm 1 Skeleton Framework
1:a language model $\mathcal{P}_{\theta}$; a prompt $P_{\tau}$; a margin $\psi$
2:a task neuron localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
3:train $\mathcal{P}_{\theta}$ via prompt tuning using $P_{\tau}$ and get $P^{\prime}_{\tau}$
4:get the validation accuracy $s_{\tau}$ for $\mathcal{P}_{\theta}$ using $P_{\tau}$
5:compute task relevance $A_{i}$ via Equation 2 and 3
6:sort target neurons in descending order by $A_{i}$
7:
8:$l\leftarrow 0$
9:$h\leftarrow 20$
10:while $l\leq r$ do                         // Binary search
11:     $m\leftarrow\lfloor(l+h)/2\rfloor$
12:     $p=m\times 0.05$
13:     get top-$p$ task neuron localized model $\mathcal{P}_{\theta_{\tau}}$
14:     train $\mathcal{P}_{\theta_{\tau}}$ via prompt tuning using $P^{\prime}_{\tau}$ and get $P^{\prime\prime}_{\tau}$
15:     get validation accuracy $s^{\prime}_{\tau}$ for $\mathcal{P}_{\theta_{\tau}}$ using $P^{\prime\prime}_{\tau}$
16:     if $s_{\tau}-s^{\prime}_{\tau}\leq\psi$ then
17:         $\mathcal{P}_{\theta_{\tau}}^{*}\leftarrow\mathcal{P}_{\theta_{\tau}}$             // The optimal model
18:         $l\leftarrow m+1$
19:     else
20:         $h\leftarrow m-1$
21:return $\mathcal{P}_{\theta_{\tau}}^{*}$
Model Method SST-2 IMDB AGNews MRPC CB
BERT (110M) Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57
Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76
Prompt-tuning 89.36 88.56 90.58 67.07 69.63
Skeleton 89.21 88.65 89.94 68.05 69.04
[…]
RoBERTa (125M) Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90
Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02
Prompt-tuning 93.84 92.72 91.87 67.23 66.06
Skeleton 92.81 91.50 90.79 67.89 69.04
[…]
Table 1: Task Neuron Localized Prompt Tuning Results. We report the accuracy of five datasets for Skeleton and other baselines. We train models using each method by three trials and report the averaged accuracy. Pruning Rate (%) and Speed up ($\times$) mean the ratio of neurons deactivated after applying our method and the inference speed up compared to the Prompt-tuning, respectively. The values in parenthesis are the number of pruned parameters. The pruning rates are calculated based on the maximum number of prunable parameters.
4.5 Details of the Skeleton Framework
Satisfying Dimension Integrity.
This section describes the detailed task localization processes; the entire process is shown in Figure 2. We prune task-irrelevant neurons from a language model. Recent language models have been trained using the transformer architecture, but they consist of complex neural computation processes. For example, sequentially connected networks are dependent on each other in terms of dimensionality. Suppose there are two sequential feed-forward networks (FFNs), $W_{1}\in\mathbb{R}^{d\times m}$ and $W_{2}\in\mathbb{R}^{m\times d}$. If we prune the first FFN and get a pruned weight $\tilde{W}_{1}\in\mathbb{R}^{d\times m^{\prime}}$, this pruning process affects the inconsistency of the dimension with the weight in the second FFN. Specifically, $W_{2}$ is forced to be pruned as $\tilde{W}_{2}\in\mathbb{R}^{m^{\prime}\times d}$. Therefore, we implement the task neuron localization algorithm by reflecting the dimension dependency of two FFNs. In addition, suppose the representation computed by the second FFN, $v\in\mathbb{R}^{d}$, is connected to the representation of a residual connection network, $u\in\mathbb{R}^{d}$. Then, if we prune the second FFN and get a pruned weight $\tilde{W}_{2}\in\mathbb{R}^{m\times d^{\prime}}$, this process also triggers the inconsistency of the dimension with the representation transferred by the residual connection, making $v\in\mathbb{R}^{d^{\prime}}$. Thus, we insert zero values into the pruned dimension of the second FFN’s representation, $v$, to ensure its integrity.
Accelerating Inference Speed.
Our framework does not just convert the pruned value of weights to zero values but eliminates the entire column (i.e., neuron) of them. Therefore, our framework accelerates the language model inference speed. We do not consider other types of networks (e.g., Attention modules, Embedding layer, Language model head) since they play a crucial role in the model’s language understanding.
Furthermore, the parameters of FFNs constitute most of the prunable parameters (about more than $65\%$); thus, pruning even only the FFNs is sufficient. The experimental results for other modules are described in Section 5.4.
5 Experiments
5.1 Experimental Setup
Datasets.
We conduct experiments on five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018). Specifically, we utilize SST-2, IMDB (sentiment analysis); AGNews (topic classification); MRPC (semantic textual matching); CB (natural language inference) to verify the applicability of our method. We split the train set and use 10% of them as a validation set.
Models and Training Methods.
We utilize two transformer variants  [^2] https://huggingface.co/docs/transformers/, BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), for our experiments since they are some of the most popular basic language models. We select a prompt tuning method (Lester et al., 2021) to compare with our method, and also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
Implementation Details.
We evaluate Skeleton and baselines on NVIDIA A5000 GPU. We train models using $l=20$ prompt tokens during 100 epochs with early-stop conditions for all datasets. In addition, we set the margin $\psi=1.0\%$ for the validation accuracy score used in the Binary search algorithm. We search $p\in[0.0,0.05,...,0.95,1.0]$ to find the best pruning rate in the Binary search algorithm. We quantify the task relevance using only $n=20$ data instances of each training dataset, following Yang et al. (2024). We specify the target modules for task neuron localization as FFN networks since the experimental results for FFN networks outperform those of other modules, shown in Figure 6. We use learning rates $\gamma_{1}\in$ [1e-5, 5e-5] and $\gamma_{2}\in$ [1e-4, 5e-3] for fine-tuning and prompt-tuning, respectively. We use a masked language model probing network including an instruction for solving a task to utilize the pre-trained knowledge of BERT and RoBERTa, as shown in Appendix A.
5.2 Skeleton is Efficient and Robust
Datasets Experiments.
We evaluate and report the task performance and efficiency of each method in Table 1 for the five natural language understanding benchmarks. We adopt BERT-base and RoBERTa-base to conduct the experiments. The results show that our method performs comparably to the prompt tuning method and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% of pruning rates). Surprisingly, the pruned networks show at most ×$\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models. We describe the detailed elapsed time (sec) of Skeleton models and prompt tuned model in Figure 3. These results reveal that our method significantly outperforms the prompt tuning method in inference speed.
Figure 3: Elapsed Time and Inference Speed-up. We plot the elapsed time and inference speed-up for BERT-base and RoBERTa-base on the four datasets. The values above the bar plot are the speed-up ratios.
Hyperparameter Experiments.
We conduct experiments to investigate the time efficiency of our framework for varying hyperparameters, as shown in Figure 4. Specifically, we search varying batch sizes and the number of input tokens to demonstrate efficiency by comparing our framework to the prompt tuning method. We sample random texts with $k$ tokens and input them into the BERT-base model for the experiments. For the batch size experiments, we evaluate the inference speed with varying batch sizes $\beta\in[16,32,64,256,512]$, fixing the number of tokens at $64$. For the number of tokens experiments, we measure the inference speed with multiple numbers of tokens $k\in[32,64,256,512]$, using the batch size of $64$. The results reveal that our framework shows robust efficiency improvement in various hyperparameters, showing better speed in higher batch sizes and a widely used number of tokens.
Figure 4: Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
5.3 Task Localization Methods Experiments
We compare our method to other task neuron selection methods: Activation, Gradient, and Random, to justify the excellence of the attribution method (Section 5.3). Activation identifies task neurons by using the activated values of neurons. We compute the sum of activated values of each neuron when inputting a dataset. Gradient selects task neurons by using the gradient values of each neuron. We compute the sum of the gradient values of each neuron when predicting the label of the utilized dataset. Random randomly selects task neurons.
We demonstrate the task localization performance of those baselines to justify our selection of the task localization method, Attribution, as shown in Figure 5. Specifically, we select the BERT-base model for the SST-2 and IMDB datasets for the experiments. We prune the top 50% of task neurons by using those methods for a fair comparison. Surprisingly, even Activation and Gradient do not show competitive performance compared to Random. However, Skeleton (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
Figure 5: Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
5.4 Module-specific Task-localization
Transformer variants have various types of modules (i.g., attention modules and FFNs). Therefore, we examine which module is suitable for maintaining the knowledge of a language model (BERT-base). Specifically, we categorize layers into four segments: (1) All layers (All), (2) Attention layers (Attn), (3) Dense layers (Dense), and (4) Feed-forward layers (FFN). All includes all linear layers. Attn consists of Query (Q), Key (K), Value (V), and Output (O) layers of the attention module. Dense contains $O$ of the attention module and two FFNs. FFN includes two FFNs. We prune the top 30% of task neurons from each module for a fair comparison. Figure 6 represents the results of our method for various types of modules. From these experiments, we reveal that FFNs are suitable for maintaining the knowledge of a language model in our prompt tuning setting. Adopting the FFN module yields faster performance than other modules without compromising accuracy.
Figure 6: Module-specific Task-localization Results. We plot the relationship between the accuracy and speed up of each module, where each diagram corresponds to the categorized module. The red-dotted line is the baseline performance (Prompt-tuning).
5.5 Prior Knowledge Injection is Important
Figure 7: Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
This section describes the importance of injecting prior task knowledge (described Section 4.4) to accurately quantify the task relevance, as shown in Figure 7. We prune the top 50% of task neurons by using two methods (i.e., Skeleton and w/o Prior Task Knowledge) for a fair comparison. The experimental results reveal that prior task knowledge injection has a crucial role in quantifying the task relevance. Excluding the Prior Knowledge Injection step results in slower convergence and lower task accuracy compared to our method.
5.6 Skeleton is Scalable to a Larger Model
We conduct experiments to verify that our method is scalable to a larger model, as shown in Table 2. Specifically, we adopt BERT-large to show the scalability of our framework. These results reveal that our method is scalable to a larger model.
Method SST-2 IMDB
BERT-base (110M) Fine-tuning (Single) 92.39 92.50
Fine-tuning (Multi) 91.36 92.16
Prompt-tuning 89.36 88.56
Skeleton 89.21 88.65
[…]
BERT-large (340M) Fine-tuning (Single) 92.77 93.57
Fine-tuning (Multi) 92.73 93.83
Prompt-tuning 91.28 89.96
Skeleton 92.19 89.35
[…]
Table 2: Experiments on various sizes of BERT.
6 Conclusion
In this study, we propose Skeleton, which explores the extreme of the prompt tuning method by localizing task-relevant neurons from the original language model. Skeleton investigates the optimal task-localized sub-network from a language model and condenses task knowledge by utilizing only task-relevant parameters during the prompt-tuning procedure. We demonstrate our method for various benchmarks and reveal that Skeleton dramatically enhances the efficiency of language models, showing significant inference speed boost-up.
Limitations
Our framework accelerates the inference speed of language models, identifying the optimal subnetwork for utilizing language models. However, while our approach exhibited comparable performance and occasionally slightly superior results to the existing prompt tuning method, these improvements were not consistent across overall task performance. In addition, our experiments are limited to five widely-used natural language understanding datasets; thus, we should extend to a wide range of datasets to validate the applicability of our framework. Furthermore, we select feed-forward networks to localize task neurons in language models since they are adequate for inference speed acceleration. However, the entire modules in Transformer should be considered as a localization target to achieve the extreme inference speed up; we leave this point as future research.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea.
Appendix A Instruction Templates.
We use text templates for input texts of the utilized datasets. Specifically, the text templates consist of an instruction and an input text, as shown in Table 3. For the prediction, we utilize the representation of the mask token (i.e., "<mask>") to utilize the pre-trained knowledge of encoder-only language models.
Appendix B Module-specific Task-localization Details
We conduct experiments for various modules in the transformer architecture, as described in Section 5.4. We implement algorithms for pruning those networks following our Skeleton framework to satisfy the dimension of adjacent networks after pruning. For example, suppose that we prune the Key network, $W^{K}\in\mathbb{R}^{d\times d}$, and get a pruned weight, $\tilde{W}^{K}\in\mathbb{R}^{d\times d^{\prime}}$, in the Attention module. Then, we should also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since they are adjacent networks. Therefore, after the pruning process, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by adding zero values to ensure the identical dimension. We also implement other layers in the same way with it.

### v3

@@
-Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning
-Abstract
-Prompt tuning methods have shown comparable performance to general training methods as parameter-efficient fine-tuning (PEFT) methods in various natural language understanding tasks.
-However, existing prompt tuning methods still utilize the entire model architecture even when solving a specific task, which prevents them from accelerating inference speed during the application procedure.
-In this paper, we propose a novel prompt tuning framework called Skeleton to efficiently utilize a language model in terms of memory and time complexity for solving various tasks, retaining only task-relevant neurons by using an explainability method.
-From our framework, we can efficiently solve various tasks by using only task-relevant neurons and prepending adequate task-specific prompt tokens with only a single language model.
-Experiments reveal that our method significantly enhances inference efficiency (at most ×$\times 1.73$ speed up) for various widely used benchmarks, showing comparable performance to the prompt tuning method.
-Moreover, our method is applicable across various transformer-based architectures, confirming its practicality and scalability.
-Skeleton: A New Framework for Accelerating Language Models via Task Neuron Localized Prompt Tuning
-Nakyeong Yang1, Jiwon Moon1, Junseok Kim2, Yunah Jang1 and Kyomin Jung1 1Seoul National University, 2Pohang University of Science and Technology {yny0506, wldnjs913, vn2209, kjung}@snu.ac.kr {junseokkim00}@postech.ac.kr
-1 Introduction
-With the success of parameter-efficient fine-tuning (PEFT) methods, pre-trained language models perform various tasks by adjusting only small numbers of parameters.
-Among PEFT methods, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) are widely used to solve multiple tasks since they require only a few tunable tokens to be prepended to the input tokens while freezing the original parameters of a language model (Chan et al., 2023; Huang et al., 2024).
-This strategy is memory efficient, as it enables a model to solve various tasks with only a few task-specific prompt tokens while retaining the original parameters.
-However, we can raise one research question regarding this task-specific trait of prompt tuning:
-"Leveraging prompt tuning, could we condense task knowledge into a few task neurons in a frozen model to efficiently utilize it?"
-Figure 1: Skeleton Framework Overview.
-Our framework activates only task-localized paths of language models in the prompt tuning procedure.
-Each color (i.g., red, blue, and yellow) corresponds to a specific task; we obtain a task-specific prompt and neuron indices after executing our framework.
+Unplug and Play Language Models: Decomposing Experts in Language Models at Inference Time
+Abstract.
+Enabled by large-scale text corpora with huge parameters, pre-trained language models operate as multi-task experts using a single model architecture.
+However, recent studies have revealed that certain neurons play disproportionately important roles in solving specific tasks, suggesting that task-relevant substructures can be isolated and selectively activated for each task.
+Therefore, we introduce Decomposition of Experts (DoE), a novel framework that dynamically identifies and activates task-specific experts within a language model to reduce inference cost without sacrificing accuracy.
+We first define a task expert as a set of parameters that significantly influence the performance of a specific task and propose a four-step unplug-and-play process: (1) receiving a user request, (2) identifying the corresponding task expert, (3) performing inference using the expert-localized model, and (4) restoring the original model and waiting for the next task.
+Using attribution methods and prompt tuning, DoE isolates task-relevant neurons, minimizing computational overhead while maintaining task performance.
+We assume a setting where a language model receives user requests from five widely used natural language understanding benchmarks, processing one task at a time.
+In this setup, we demonstrate that DoE achieves up to a $\times 1.73$ inference speed-up with a 65% pruning rate, without compromising accuracy.
+Comparisons with various task expert localization methods reveal that DoE effectively identifies task experts, while ablation studies validate the importance of its components.
+Additionally, we analyze the effects of batch size, token count, and layer types on inference speed-up, providing practical insights for adopting DoE.
+The proposed framework is both practical and scalable, applicable to any transformer-based architecture, offering a robust solution for efficient task-specific inference.
+Figure 1.
+Decompostion of Experts (DoE) Framework Overview.
+Our framework leverages an unplug-and-play strategy by identifying and activating only the task expert of a language model during the inference procedure when given a user request.
+Each color (i.e., red, blue, and yellow) corresponds to a specific task; we utilize a task-specific prompt and neuron indices obtained through a prompt tuning process.
 Gray-colored paths in the figure represent task-irrelevant paths that are never used during training and inference, thereby improving inference speed.
-Existing prompt tuning frameworks are limited by the need to use the full original model architecture during inference.
-In fact, these methods may require additional inference time, as the prepended prompt tokens introduce further computational overhead.
+Knowledge Localization, Decomposition of Experts, Language Model
+journalyear: 2025copyright: noneconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/3746252.3761011ccs: Computing methodologies Natural language processing
+1. Introduction
+Training on large-scale text corpora with numerous parameters enables recent pre-trained language models to perform a wide range of tasks using a single model architecture (Devlin et al., 2018; Liu et al., 2019; Lewis et al., 2019; Touvron et al., 2023; OpenAI, 2023).
+In this context, previous studies have supposed that each neuron in a language model may have a distinct role; thus, they have tried to investigate the role of individual neurons to explain task-specific knowledge in language models (Panigrahi et al., 2023; Yang et al., 2024; Chen et al., 2025).
+They have addressed various tasks, including sentiment analysis, natural language inference, and social bias detection.
+Therefore, this line of studies has laid the groundwork for identifying task-relevant neurons in language models by validating that these neurons possess sufficient knowledge to solve a specific task.
+Based on these prior studies, we propose a novel framework, Decomposition of Experts (DoE), that improves the efficiency of language models by dynamically selecting task-specific experts in response to user requests.
+This approach reduces inference time in single-task scenarios.
+We first define a task expert as a subset of parameters that significantly contribute to solving a specific task.
+Based on this definition, we describe the DoE framework as an unplug-and-play language model composed of four stages: (1) receiving a user request, (2) identifying the corresponding task expert, (3) performing inference using the expert-localized model, and (4) restoring the original model and waiting for the next task (see Figure 1).
+Specifically, we employ the attribution method (Yang et al., 2024) to identify and activate only task-relevant neurons, which we treat as the task expert.
+DoE also adopts prompt tuning (Lester et al., 2021) to align task knowledge more effectively within parameters, enabling expert selection using a minimal number of neurons.
+While previous pruning approaches have primarily focused on permanently removing unnecessary neurons or connections to reduce the overall parameter count, they typically rely on fixed subnetwork architectures optimized for specific tasks or consistently utilize lightweight models during inference (Frankle and Carbin, 2018; Ma et al., 2023; Chen et al., 2024).
+In contrast, our method assumes that pre-trained language models inherently possess sufficient multi-task capability.
+Rather than statically removing parameters, we achieve efficiency by dynamically selecting task-specific expert neurons and temporarily deactivating irrelevant ones (unplug).
+These neurons can then be reactivated for different tasks, enabling a reversible and task-adaptive computation flow.
+From this perspective, our framework offers a novel efficiency strategy that lies at the intersection of pruning and modularization.
+It is particularly advantageous in multi-task settings where rapid task switching is essential.
+DoE dramatically accelerates the inference speed of language models since it uses only a task expert, which is a small part of the entire model during inference.
+For example, to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only a task expert to resolve the task.
+Our experimental settings assume that a language model is given user requests about five widely used natural language understanding benchmarks.
+The experimental results demonstrate that our method significantly enhances inference efficiency without sacrificing the original performance of each task.
+Surprisingly, our method improves the inference speed by up to $\times 1.73$ with a 65% parameter pruning rate while solving a specific task.
+In addition, we conduct experiments to compare various task expert selection methods and reveal that the attribution method effectively identifies task experts.
+We also investigate the effects of batch size, the number of tokens, and types of layers on inference speed-up.
+Our framework is practical and scalable, as it can be applied to any transformer-based architecture.
+In summary, this work makes the following contributions:
+  • We first define task expert and propose Decomposition of Expert (DoE), a novel language model usage framework that accelerates inference speed by identifying and isolating a task expert within a language model, leveraging an unplug-and-play strategy.
+  • We demonstrate that DoE effectively identifies task experts over other task expert localization methods and enhances the inference speed of each task without compromising task performance.
+  • We offer various practical guidelines for adopting DoE by investigating the effects of hyper-parameters on inference speed-up and evaluating the scalability of the method.
+2. Backgrounds
+2.1. Task Knowledge Localization
+Prior works have shown that language models trained with a vast size of corpora with a huge number of parameters can serve as a multi-task expert (Radford et al., 2019; Raffel et al., 2020).
+These multi-tasking abilities are attributed to exploiting commonalities and differences recognized while training multiple categories of texts (Zhang et al., 2022; Chen et al., 2024).
+Recent studies have shown that multi-task knowledge is distributed across sparse neurons in language models (Liu et al., 2023; Chan et al., 2023).
+Building on this observation, earlier research has investigated the roles of individual parameters in handling specific tasks (Panigrahi et al., 2023; Chen et al., 2025; Yang et al., 2024).
+Panigrahi et al.
+(2023) have suggested a training-based task knowledge localization method, model grafting, which identifies task-relevant neurons by training new parameters to mask original parameters.
+Chen et al.
+(2025) have investigated the query-relevant neurons by utilizing intergrated gradients (Sundararajan et al., 2017), an explainability method that derives the importance of each neuron when inferring a long text.
+Yang et al.
+(2024) have illuminated bias-relevant neurons by utilizing attribution (Shrikumar et al., 2016), an explainability method that derives the importance of each neuron when inferring a specific biased text.
+Yang et al.
+(2024) have verified that the attribution effectively detects bias-relevant neurons and proposed a method applicable to language modeling.
+Although Panigrahi et al.
+(2023); Chen et al.
+(2025); Yang et al.
+(2024) have effectively identified task-relevant neurons, they have not proposed practical methods to enhance the usability of language models.
+Panigrahi et al.
+(2023) requires excessive additional storage and computation for newly trained masking parameters, equivalent to the number of model parameters.
+In addition, Yang et al.
+(2024) has a limitation in just eliminating the knowledge-relevant neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded.
+In other words, this line of studies has only focused on investigating knowledge-relevant neurons, failing to propose a practical method for efficiently utilizing language models during the inference procedure.
+2.2. Language Model Efficiency
+Existing studies have proposed various approaches to enhance the efficiency of language models, which can be categorized into three main groups: (1) knowledge distillation, (2) quantization, and (3) pruning.
+Knowledge distillation transfers the knowledge of larger or stronger models into smaller or more efficient ones (Gu et al., 2023; Agarwal et al., 2024).
+Quantization reduces the precision of model parameters and activations to lower memory and computational costs (Shao et al., 2023; Yue et al., 2024; Xu et al., 2024).
+Pruning eliminates less important parameters to reduce model size and inference time (Frankle and Carbin, 2018; Ma et al., 2023; Chen et al., 2024).
+While these methods have improved efficiency, they typically rely on statically modifying the model into fixed architectures tailored for specific tasks and deploying task-specific lightweight models consistently at inference time.
+In contrast, our approach proposes a new framework of language model usage: Unplug-and-plug language models: preserving the full model and dynamically activating only task-relevant neurons, enabling adaptive and reversible task-specific inference without architectural reconfiguration.
+2.3. Parameter-efficient Fine-tuning.
+Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023).
+Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various natural language processing tasks (Wang et al., 2023; Bansal et al., 2023).
+Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model.
+However, they are still limited as they are not parameter-efficient during the inference process.
 Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) have attempted to accelerate the inference speed of prompt-tuned models by pruning unnecessary prompt tokens.
 However, they have only marginally improved inference efficiency, as the number of pruned tokens is a small fraction ($\leq 0.1\%$) of all the parameters of a language model (Lester et al., 2021).
-Prior studies have attempted to investigate and analyze task-relevant neurons in language models to explain the language models’ behavior (Panigrahi et al., 2023; Yang et al., 2024).
-Specifically, Panigrahi et al.
-(2023) has suggested training new parameters, which mask the original parameters of language models to determine task neurons.
-Yang et al.
-(2024) has utilized an explainability method to determine task-relevant neurons and reveal that some task-relevant neurons significantly affect the outputs of specific tasks (i.e., natural language understanding, social bias).
-However, while they have introduced a method to identify task-relevant neurons, they fall short of proposing a practical approach to efficiently utilize those neurons during training and inference.
-To address this challenge, we propose a novel training framework, Skeleton, designed to enhance the efficiency of language models by leveraging only task-relevant neurons for solving various tasks.
-Specifically, we use attribution (Yang et al., 2023, 2024), an explainability method, to identify and utilize only task-relevant neurons in the prompt tuning procedure.
-Our framework requires only a few prompt tokens to be updated and freezes the original parameters of a language model, as it adopts the prompt tuning method.
-Furthermore, Skeleton dramatically accelerates the inference speed of language models since it utilizes only task-relevant neurons.
-For example, in order to solve a specific task (e.g., sentiment analysis, natural language inference), our method prepends prompt tokens related to the task and activates only the task-relevant neurons to solve the task.
-We demonstrate our method in various widely used natural language understanding benchmarks and reveal that our method significantly outperforms existing PEFT methods in inference efficiency without sacrificing the original performance of each task.
-Surprisingly, our method improves the inference speed up to ×$\times 1.73$ with 65% of the parameters pruning rate.
-We additionally conduct various analyses to offer a guideline for using our new prompt tuning framework, such as investigating which types of layers are crucial for processing task knowledge in a language model.
-Furthermore, our prompt tuning framework is practical and scalable since it is applicable to any transformer-based architecture.
-In summary, this work makes the following contributions:
-  • We first define task neuron and propose Task Neuron Localized Prompt Tuning (Skeleton), a practical training framework for accelerating the inference speed of a language model for various tasks.
-  • We reveal that Skeleton successfully enhances the inference speed of each task without compromising the task performance.
-2 Backgrounds
-2.1 Parameter-efficient Fine-tuning
-Parameter-efficient Fine-tuning (PEFT) methods have emerged as a standard of efficient training methods for language models (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021; Hu et al., 2021; Poth et al., 2023).
-Among them, prompt tuning methods (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021) have been adopted for various NLP tasks (Wang et al., 2023; Bansal et al., 2023).
-Prompt tuning methods only update prompt tokens while maintaining the original parameters of a language model; thus, this characteristic makes the training process memory-efficient and effective in injecting task knowledge into a language model.
-However, they still face a limitation as they are not parameter-efficient during the inference process.
-Therefore, existing methods (Ma et al., 2022; Liang et al., 2023) propose prompt token selection methods to enhance the inference speed of prompt-tuned language models.
-Nevertheless, they have only marginally enhanced inference efficiency since the number of pruned tokens is only a tiny proportion of the entire number of input tokens.
-2.2 Task Knowledge Localization
-In spite of the significant performance of language models, it remains challenging to accurately investigate the role of each parameter in language models in dealing with a specific task.
-Therefore, existing studies have attempted to identify task-specific neurons in executing a specific task (Panigrahi et al., 2023; Yang et al., 2023, 2024).
-Panigrahi et al.
-(2023); Yang et al.
-(2023, 2024) have studied to localize neurons for performing a specific task by clarifying the role of each parameter in a language model.
-Panigrahi et al.
-(2023) has suggested a training-based task knowledge localization method, model grafting, which localizes a task knowledge by training new parameters to mask original parameters.
-Although it has effectively identified task neurons, it has required excessive additional storage for newly trained masking parameters, which is equivalent to the number of model parameters.
-Yang et al.
-(2023, 2024) has illuminated task neurons by utilizing the attribution technique (Shrikumar et al., 2016), an explainability method that derives the importance of each feature when solving a specific task.
-Yang et al.
-(2023, 2024) has verified that the attribution effectively detects task neurons for solving a specific task and proposed a task neuron detection method applicable to language modeling tasks.
-However, Yang et al.
-(2023, 2024) has a limitation in just pruning the task neurons; if we prune a large number of neurons, the performance of a specific task can be significantly degraded.
-Furthermore, the prior study has only proposed the method for pruning task neurons and is far from proposing a practical method to improve the inference speed of task-localized language models.
-The architecture of transformer variants consists of complex neural computation processes; thus, satisfying the integrity of dimension dependency for adjacent networks while pruning the weights is challenging.
-We adopt the attribution method (Yang et al., 2023, 2024) and propose a new practical framework for accelerating a language model, where the framework utilizes only task neurons while training a language model via prompt tuning, resolving the challenges in pruning networks in transformer variants.
-3 Problem Definition
-3.1 Prompt Tuning
-We adopt the prompt tuning method to propose a new framework for accelerating a language model.
-Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)=\prod_{k=1}^{K}\mathcal{P}_{\theta}(y_{k}|x,y_{1},...,y_{k-1})$ that represents a language model, where $\theta$ is parameters of the language model.
+Our DoE framework adopts the prompt tuning method (Lester et al., 2021) to align the original task knowledge when localizing a task expert.
+3. Problem Definition
+3.1. Unplug-and-play language model
+We define unplug-and-play language model as a language model that satisfies the following two conditions: (1) it should perform a single task using only a tiny fraction of its parameters, and (2) the model architecture is fully reversible; thus, it can restore its original parameters after task localization, allowing the model to remain idle to be available for other tasks.
+Typical compression methods, such as pruning and knowledge distillation, cannot be considered unplug-and-play language models since they statically modify the model into fixed architectures for specific tasks and consistently deploy task-specific lightweight models during inference (irreversible).
+3.2. Prompt Tuning
+We adopt the prompt tuning method to propose a new framework for efficiently utilizing a language model.
+Formally, suppose we have a function $\mathcal{P}_{\theta}(y|x)$ that represents a language model, where $\theta$ is the parameters of the language model.
 Prompting is the method of adding extra information to the model to condition during its generation of $y$.
@@
 After the prompt tuning, we get the new conditional generation function $\mathcal{P}_{\theta}(y|[P^{\prime}_{\tau};x])$.
-3.2 Skeleton Framework
-In the Skeleton framework, we still update only prompt tokens, freezing the original parameters.
-However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we first define task neuron as follows:
-Task Neuron.
+However, we assume that there is a minimal parameter combination $\theta_{\tau}\subset\theta$ for solving a specific task $\tau$; thus, we define task expert as follows:
+3.3. Task Expert
 Let $\mathcal{P}_{\theta}$ be a language model, where $\theta=\{\theta_{1},...,\theta_{N}\}$ is a set of parameters of it.
-Then, $\theta_{\tau}\subset\theta$ is defined as task neurons for a task $\tau$ if the below equation is satisfied:
-$\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta}([P^{\prime}_{\tau};x]))\approx\sum_{(x,y)\in\mathcal{D}_{\tau}}\mathcal{L}(y,\mathcal{P}_{\theta_{\tau}}([P^{\prime\prime}_{\tau};x]))$ (1)
-where $\mathcal{D}_{\tau}$ means a dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy).
-$P^{\prime}_{\tau}$ and $P^{\prime\prime}_{\tau}$ are task-specific prompts.
-$N$ means the total number of parameters in a language model.
-Skeleton framework adopts the prompt tuning strategy using only task neurons $\theta_{\tau}$, and get the new task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime\prime}_{\tau};x])$.
-Therefore, we can utilize the language model more time-efficiently by using only those neurons within the Skeleton framework.
-In addition to being time efficient, this framework is also memory-efficient since it requires only a small fraction of original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime\prime}_{\tau}$.
-4 Methods
-In this section, we describe the process of Skeleton.
-We first train a language model using the prompt tuning method to inject prior knowledge of a specific task into it.
-After that, we quantify the task relevance of each neuron to determine task neurons in a language model.
-We finally retrain the prompt tokens with the prompt tuning method to condense the task knowledge to only the task neurons.
-Note that the original language model maintains only task neurons via a pruning method in this process.
-After applying Skeleton framework, we can utilize the language model efficiently by using only the task-specific prompt and task neurons, accelerating the inference speed for a specific task.
-The entire process of Skeleton is described in Figure 2 and Algorithm 1.
-Figure 2: Task Neuron Localization Process of Skeleton.
-After training a model using our framework, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure.
+Then, $\theta_{\tau}\subset\theta$ is called as a task expert for a task $\tau$ if the below formula is satisfied:
+(1) $\scriptsize\sum_{(x,y)\in\mathcal{D}_{\tau}}\hskip-9.10509pt\mathcal{L}(\mathcal{P}_{\theta},x,y)=\epsilon+\hskip-8.5359pt\sum_{(x,y)\in\mathcal{D}_{\tau}}\hskip-9.10509pt\mathcal{L}(\mathcal{P}_{\theta_{\tau}},x,y)\small$
+where $\mathcal{D}_{\tau}$ denotes the dataset for a task $\tau$ and $\mathcal{L}$ is a score function (e.g., loss function or accuracy) evaluated on the given model.
+$N$ means the total number of parameters in the language model.
+$\epsilon$ is a small error value.
+In other words, if the scores computed on $\mathcal{P}_{\theta}$ and $\mathcal{P}_{\theta_{\tau}}$ for the given dataset $\mathcal{D}_{\tau}$ are approximately the same, we can utilize only the task neurons $\theta_{\tau}$ to solve the task $\tau$.
+DoE framework adopts the prompt tuning strategy using only the task expert $\theta_{\tau}$ in a language model; thus, we get the task-localized conditional generation function $\mathcal{P}_{\theta_{\tau}}(y|[P^{\prime}_{\tau};x])$ instead of $\mathcal{P}_{\theta_{\tau}}(y|x)$.
+Notice that since the prompt tuning strategy does not modify the original parameters of language models, a model that adopts the prompt tuning can still be used as a unplug-and-play language model.
+We can utilize the language model more time-efficiently by using only the sparse neurons within our  framework.
+In addition to being time-efficient, this framework is also memory-efficient since it requires only a small fraction of the original parameters $\theta_{\tau}$ and just a few task-specific prompt tokens $P^{\prime}_{\tau}$.
+4. Decomposition of Expert (DoE)
+In this section, we describe the process for adopting the DoE framework.
+We first train a language model using the prompt tuning method to align the prior knowledge of a specific task.
+After that, we quantify the task relevance of each neuron to determine a task expert in a language model.
+Finally, we apply prompt tuning to retrain the prompt tokens, enabling the condensation of task knowledge into the task expert.
+During this process, only task-relevant neurons are preserved in the original language model.
+The method is performed for a set of predefined $T$ tasks; thus, we have $T$ task-specific prompts and $T$ sets of task experts.
+After applying the DoE framework, we can utilize the language model efficiently by using only the task-specific prompt and task expert, accelerating the inference speed for a specific task.
+The entire process of DoE is described in Figure 2 and Algorithm 1.
+Figure 2.
+Task Expert Localization Process of DoE.
+After training a model with the prompt tuning, a task prompt and neurons are assigned to each task, and we utilize only those task-specific resources in the inference procedure.
 Red-colored values in the prune weights, $\tilde{W}_{1}$ and $\tilde{W}_{2}$, mean task neurons for a specific task (task A).
 The weight of FFN2 is pruned depending on FFN1’s dimension, and the output representation of FFN2 is zero-recovered (Blue-colored values) to match the output dimension of the residual network (described in Section 4.5).
-4.1 Quantifying Task Relevance of Neurons
+4.1. Task Relevance Quantification
 This section describes the process of computing the task relevance of each neuron.
-We adopt an attribution method (Yang et al., 2023, 2024) to quantify the skill relevance of neurons in the pre-trained language models.
+We adopt an attribution method (Yang et al., 2024) to quantify the skill relevance of neurons in pre-trained language models.
 Formally, given a task-specific prompt tokens $P_{\tau}$ and text input tokens $x$ provided to $\mathcal{P}_{\theta}$, the contribution of the $i$-th neuron to the prediction of an output $y$ is defined as follows:
-$A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$ (2)
-where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])/\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron.
+(2) $A^{(P_{\tau},y,x)}_{i}(h)=|h_{i}\times\frac{\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])}{\partial h_{i}}|$
+where $h$ corresponds to a hidden representation, and $\partial\mathcal{P}_{\theta}(y|[P_{\tau};x])$ $\partial h_{i}$ means the gradient of the logit $\mathcal{P}_{\theta}(y|[P_{\tau};x])$ with respect to the $i$-th neuron of the representation $h$.
 We adopt transformer variants; thus, activations and gradients are computed for all input token representations.
@@
 In addition, there are multiple data instances for each task $\tau$; thus, we aggregate attributions for tokens and instances as follows:
-$A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)=\sum_{j}^{N}\sum_{k}^{K_{j}}\frac{1}{K_{j}}A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ (3)
+(3) $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)=\sum_{j}^{N}\sum_{k}^{K_{j}}\frac{1}{K_{j}}A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$
 where $A^{(P_{\tau},y_{j},x_{j},t_{k})}_{i}(h)$ means the attribution scores computed for an input text token $t_{k}\in x_{j}$.
@@
 To simplify the notation, $A^{(P_{\tau},\mathcal{D}^{\tau})}_{i}(h)$ will be abbreviated as $A_{i}$ from this point forward.
-4.2 Identifying Task-relevant Neurons.
+4.2. Task Expert Identification
 After quantifying each neuron’s task relevance, we determine task-relevant neurons and eliminate task-irrelevant neurons using a structured pruning method.
@@
 We define a margin $\psi$ and regard $\theta_{\tau}$ as task neurons if the accuracy degradation after neuron localization is smaller than the specified margin $\psi$.
-4.3 Task Neuron Localization.
+4.3. Task Expert Localization
 After identifying task neurons, we eliminate task-irrelevant neurons from a language model using a structured pruning method.
@@
 After localizing a language model, we further intensify task knowledge of a localized language model using a prompt tuning method.
-4.4 Task Neuron Localized Prompt Tuning
-Our Skeleton framework follows the four steps to train a model: (1) the first training step for injecting prior task knowledge; for injecting prior task knowledge; (2) the quantification step of computing task relevance for each neuron; (3) the task neuron localization step; (4) the final prompt tuning step for condensing task knowledge.
-The Prior Knowledge Injection Step.
-The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a specific task in order to accurately quantify the task-specific knowledge of each neuron.
-Therefore, we first execute the prior knowledge injection process.
-Specifically, we train a language model $\mathcal{P}_{\theta}$ via prompt tuning and get a task-specific prompt $P_{\tau}^{\prime}$.
-The Quantification Step.
+4.4. The Entire Process of DoE
+Our DoE framework follows the three steps: (1) the quantification step of computing task relevance for each neuron; (2) the task expert determination and localization step; (3) the final prompt tuning step for condensing task knowledge.
+The algorithm for the entire process is described in Algorithm 1.
+(1) The Knowledge Quantification Step.
+The attribution method quantifies the contribution of each neuron to the prediction of an output; thus, a language model must have the ability to solve a task in order to quantify the task-specific knowledge of each neuron accurately.
+Therefore, we first execute the prompt tuning to align the knowledge in the language model.
+Specifically, we conduct the prompt tuning using a frozen language model $\mathcal{P}_{\theta}$ and get an updated task-specific prompt $P_{\tau}^{\prime}$.
 After injecting task knowledge into the language model, we quantify the task relevance $A_{i}$ of each neuron for $\mathcal{P}_{\theta}$ and $P_{\tau}^{\prime}$ using Equation 2 and 3.
-The Task Neuron Localization Step.
+(2) The Task Expert Localization Step.
 After quantifying the task relevance $A_{i}$, we first sort the neurons by the task relevance in descending order.
@@
 Note that we use a Binary search algorithm to find the optimal ratio of neurons $p\in[0.0,0.05,...,0.95,1.0]$.
-This search process requires only a maximum of five trials since we search pruning ratio in 0.05 units.
-The Knowledge Condensation Step.
-During the task neuron localization step, we condense the task knowledge into only the task neurons.
-Although we never update the original parameters of the language model, we can condense the task knowledge into the task neurons by updating only $P_{\tau}^{\prime}$.
-We train the task neuron localized model $P_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$.
+This search process requires at most five trials, as the pruning ratio is explored in increments of 0.05.
+(3) The Knowledge Condensation Step.
+During the task expert localization step, we condense the task knowledge to only the task expert.
+Although we never update the original parameters of the language model, we can condense the task knowledge into the task expert by updating only $P_{\tau}^{\prime}$.
+We train the task neuron localized model $\mathcal{P}_{\theta_{\tau}}$ and get a new knowledge-condensed task prompt $P_{\tau}^{\prime\prime}$.
 The Binary search algorithm determines the optimal pruning rate after this step.
-Algorithm 1 Skeleton Framework
+Algorithm 1 DoE Framework
 1:a language model $\mathcal{P}_{\theta}$; a prompt $P_{\tau}$; a margin $\psi$
-2:a task neuron localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
-3:train $\mathcal{P}_{\theta}$ via prompt tuning using $P_{\tau}$ and get $P^{\prime}_{\tau}$
+2:a task-localized model $\mathcal{P}^{*}_{\theta_{\tau}}$
+3:update $P_{\tau}$ via prompt tuning using $\mathcal{P}_{\theta}$ and get $P^{\prime}_{\tau}$
 4:get the validation accuracy $s_{\tau}$ for $\mathcal{P}_{\theta}$ using $P_{\tau}$
 5:compute task relevance $A_{i}$ via Equation 2 and 3
-6:sort target neurons in descending order by $A_{i}$
+6:sort neurons in descending order by $A_{i}$
 7:
-8:$l\leftarrow 0$
-9:$h\leftarrow 20$
+[…]
+[…]
 10:while $l\leq r$ do // Binary search
@@
 12:  $p=m\times 0.05$
-13:     get top-$p$ task neuron localized model $\mathcal{P}_{\theta_{\tau}}$
-14:     train $\mathcal{P}_{\theta_{\tau}}$ via prompt tuning using $P^{\prime}_{\tau}$ and get $P^{\prime\prime}_{\tau}$
+13:  get top-$p$ task-localized model $\mathcal{P}_{\theta_{\tau}}$
+14:  update $P^{\prime}_{\tau}$ via prompt tuning using $\mathcal{P}_{\theta_{\tau}}$ and get $P^{\prime\prime}_{\tau}$
 15:  get validation accuracy $s^{\prime}_{\tau}$ for $\mathcal{P}_{\theta_{\tau}}$ using $P^{\prime\prime}_{\tau}$
@@
 20:    $h\leftarrow m-1$
-21:return $\mathcal{P}_{\theta_{\tau}}^{*}$
-Model Method SST-2 IMDB AGNews MRPC CB
-BERT (110M) Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57
-Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76
-Prompt-tuning 89.36 88.56 90.58 67.07 69.63
-Skeleton 89.21 88.65 89.94 68.05 69.04
-[…]
-[…]
-RoBERTa (125M) Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90
-Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02
-Prompt-tuning 93.84 92.72 91.87 67.23 66.06
-Skeleton 92.81 91.50 90.79 67.89 69.04
-[…]
-[…]
-Table 1: Task Neuron Localized Prompt Tuning Results.
-We report the accuracy of five datasets for Skeleton and other baselines.
-We train models using each method by three trials and report the averaged accuracy.
-Pruning Rate (%) and Speed up ($\times$) mean the ratio of neurons deactivated after applying our method and the inference speed up compared to the Prompt-tuning, respectively.
-The values in parenthesis are the number of pruned parameters.
-The pruning rates are calculated based on the maximum number of prunable parameters.
-4.5 Details of the Skeleton Framework
+21:  end if
+22:end while
+23:return $\mathcal{P}_{\theta_{\tau}}^{*}$
+4.5. Details of the DoE Framework
 Satisfying Dimension Integrity.
-This section describes the detailed task localization processes; the entire process is shown in Figure 2.
+This section describes the detailed task expert localization processes; the entire process is shown in Figure 2.
 We prune task-irrelevant neurons from a language model.
@@
 Specifically, $W_{2}$ is forced to be pruned as $\tilde{W}_{2}\in\mathbb{R}^{m^{\prime}\times d}$.
-Therefore, we implement the task neuron localization algorithm by reflecting the dimension dependency of two FFNs.
+Therefore, we implement the task expert localization algorithm by reflecting the dimension dependency of two FFNs.
 In addition, suppose the representation computed by the second FFN, $v\in\mathbb{R}^{d}$, is connected to the representation of a residual connection network, $u\in\mathbb{R}^{d}$.
@@
 Thus, we insert zero values into the pruned dimension of the second FFN’s representation, $v$, to ensure its integrity.
+If we do not insert zero values, we can not fully utilize the knowledge of the residual connection.
 Accelerating Inference Speed.
@@
 We do not consider other types of networks (e.g., Attention modules, Embedding layer, Language model head) since they play a crucial role in the model’s language understanding.
-Furthermore, the parameters of FFNs constitute most of the prunable parameters (about more than $65\%$); thus, pruning even only the FFNs is sufficient.
+Furthermore, the parameters of FFNs constitute most of the prunable parameters (about $70\%$); thus, pruning even only the FFNs is sufficient.
 The experimental results for other modules are described in Section 5.4.
-5 Experiments
-5.1 Experimental Setup
+5. Experiments
+5.1. Experimental Setup
 Datasets.
-We conduct experiments on five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018).
+We suppose that language models are given requests about five widely-used natural language understanding datasets [^1] https://huggingface.co/docs/datasets/ (Maas et al., 2011; Zhang et al., 2015; Wang et al., 2018) from users.
 Specifically, we utilize SST-2, IMDB (sentiment analysis); AGNews (topic classification); MRPC (semantic textual matching); CB (natural language inference) to verify the applicability of our method.
-We split the train set and use 10% of them as a validation set.
+We split the train sets and use 10% of them as a validation set.
+Model Method SST-2 IMDB AGNews MRPC CB Avg
+BERT
+Fine-tuning (Single) 92.39 92.50 94.12 82.35 78.57 87.98
+Fine-tuning (Multi) 91.36 92.16 93.98 78.02 79.76 87.05
+Prompt-tuning 89.36 88.56 90.58 67.07 69.63 81.04
+DoE (ours) 89.21 88.65 89.94 68.05 69.04 80.98
+[…]
+[…]
+RoBERTa
+Fine-tuning (Single) 94.34 94.56 94.61 85.13 86.90 91.10
+Fine-tuning (Multi) 94.65 94.70 94.64 86.85 72.02 88.57
+Prompt-tuning 93.84 92.72 91.87 67.23 66.06 82.34
+DoE (ours) 92.81 91.50 90.79 67.89 69.04 82.40
+[…]
+[…]
+Table 1.
+Decomposition of Experts Results.
+We report the accuracy of five datasets for DoE and other baselines, such as two fine-tuning methods (upper bounds) and prompt-tuning.
+We train models using each method by three trials and report the averaged accuracy.
+Our method, DoE, achieves higher speed-ups (up to $\times$1.73) than other methods by leveraging high pruning rates (up to 65%), while maintaining comparable performance to the existing prompt tuning method.
 Models and Training Methods.
 We utilize two transformer variants [^2] https://huggingface.co/docs/transformers/, BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), for our experiments since they are some of the most popular basic language models.
-We select a prompt tuning method (Lester et al., 2021) to compare with our method, and also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
+We select a prompt tuning method (Lester et al., 2021) as the backbone and baseline of our method since it is one of the basic PEFT methods.
+We note that the typical LoRA (Hu et al., 2021) method is not appropriate for DoE, as it adds low-rank updates to the original weights at inference time, making parameter pruning difficult.
+Additionally, we also include the results of basic fine-tuning and multi-task fine-tuning to show the upper-bound performance of each task.
 Implementation Details.
-We evaluate Skeleton and baselines on NVIDIA A5000 GPU.
+We evaluate DoE and baselines on NVIDIA A5000 GPU.
 We train models using $l=20$ prompt tokens during 100 epochs with early-stop conditions for all datasets.
-In addition, we set the margin $\psi=1.0\%$ for the validation accuracy score used in the Binary search algorithm.
+In addition, we set the margin $\psi=1.0\%p$ for the validation accuracy score used in the Binary search algorithm.
 We search $p\in[0.0,0.05,...,0.95,1.0]$ to find the best pruning rate in the Binary search algorithm.
@@
 (2024).
-We specify the target modules for task neuron localization as FFN networks since the experimental results for FFN networks outperform those of other modules, shown in Figure 6.
+We specify the target modules for task expert localization as FFN networks since the experimental results for FFN networks outperform those of other modules (Described in Section 5.4).
 We use learning rates $\gamma_{1}\in$ [1e-5, 5e-5] and $\gamma_{2}\in$ [1e-4, 5e-3] for fine-tuning and prompt-tuning, respectively.
-We use a masked language model probing network including an instruction for solving a task to utilize the pre-trained knowledge of BERT and RoBERTa, as shown in Appendix A.
-5.2 Skeleton is Efficient and Robust
-Datasets Experiments.
-We evaluate and report the task performance and efficiency of each method in Table 1 for the five natural language understanding benchmarks.
+We predict output texts using a masked language model probing network for the prompts that include an instruction for solving a task, as shown in Table 3.
+Specifically, the text templates consist of an instruction and an input text.
+For the prediction, we use the representation of the mask token (i.e., “<<mask>>”).
+5.2. DoE is Efficient and Robust
+Unplug-and-play Experiments.
+We evaluate and report the task performance and efficiency of the baselines in Table 1 for the five natural language understanding benchmarks.
 We adopt BERT-base and RoBERTa-base to conduct the experiments.
-The results show that our method performs comparably to the prompt tuning method and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% of pruning rates).
-Surprisingly, the pruned networks show at most ×$\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models.
-We describe the detailed elapsed time (sec) of Skeleton models and prompt tuned model in Figure 3.
-These results reveal that our method significantly outperforms the prompt tuning method in inference speed.
-Figure 3: Elapsed Time and Inference Speed-up.
+The results show that our method performs comparably to the prompt tuning method, which utilizes the full model architecture, and even outperforms it in accuracy for some datasets using only a small number of task-localized parameters (e.g., at most 65% pruning rates).
+Surprisingly, the pruned networks show at most $\times 1.72$ speed up; thus, these results reveal that our method successfully localizes efficient task-relevant sub-networks from language models.
+We describe the detailed elapsed time (sec) of DoE models and prompt tuned model in Figure 3.
+These reveal that our method significantly outperforms the prompt tuning method in inference speed.
+Figure 3.
+Elapsed Time and Inference Speed-up.
 We plot the elapsed time and inference speed-up for BERT-base and RoBERTa-base on the four datasets.
@@
 For the number of tokens experiments, we measure the inference speed with multiple numbers of tokens $k\in[32,64,256,512]$, using the batch size of $64$.
+We conduct experiments on the fixed pruning rates $30\%$ and $50\%$ for the comparison.
 The results reveal that our framework shows robust efficiency improvement in various hyperparameters, showing better speed in higher batch sizes and a widely used number of tokens.
-Figure 4: Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
-5.3 Task Localization Methods Experiments
+Figure 4. Inference Speed Experiments for Varying Hyper-parameters (Batch size & # of Tokens). We report the results for the pruning rates of 30% and 50% for BERT-base to ensure a fair comparison.
+Experiments on various sizes of BERT.
+We conduct experiments to evaluate the scalability of our method to larger models, as shown in Table 2.
+Specifically, we adopt BERT-large to show the scalability of our framework.
+These results reveal the possibility that our method can be scaled to a larger model.
+Method SST-2 IMDB
+BERT-base Fine-tuning (Single) 92.39 92.50
+Fine-tuning (Multi) 91.36 92.16
+Prompt-tuning 89.36 88.56
+DoE 89.21 88.65
+[…]
+[…]
+BERT-large Fine-tuning (Single) 92.77 93.57
+Fine-tuning (Multi) 92.73 93.83
+Prompt-tuning 91.28 89.96
+DoE 92.19 89.35
+[…]
+[…]
+Table 2. Experiments on various sizes of BERT.
+5.3. Task Localization Methods Comparison
 We compare our method to other task neuron selection methods: Activation, Gradient, and Random, to justify the excellence of the attribution method (Section 5.3).
@@
 Surprisingly, even Activation and Gradient do not show competitive performance compared to Random.
-However, Skeleton (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
-Figure 5: Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
-5.4 Module-specific Task-localization
-Transformer variants have various types of modules (i.g., attention modules and FFNs).
+However, DoE (i.e., Attribution) outperforms other methods in the two datasets, revealing the excellence of our method.
+Figure 5. Task Localization Methods Experiments. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
+Dataset Template Label
+SST-2 Text: {text}. The sentiment of the text is <<mask>>. positive / negative
+IMDB Text: {text}. The sentiment of the text is <<mask>>. positive / negative
+AGNews Text: {text}. The topic of the text is <<mask>>. World / Sports / Business / Science
+MRPC Text1: {text1}. Text2: {text2}. The two texts are <<mask>>. different / equivalent
+CB Premise: {premise}. Hypothesis: {hypothesis}. The premise and hypothesis have a relationship of <<mask>>. implication / contradiction / neutrality
+Table 3. Instruction Templates and Labels for Each Dataset. The templates show how text, premise, hypothesis, and mask tokens are used in each dataset.
+5.4. Module-specific Task-localization
+Transformer variants have various types of modules (e.g., attention modules and FFNs).
 Therefore, we examine which module is suitable for maintaining the knowledge of a language model (BERT-base).
@@
 Adopting the FFN module yields faster performance than other modules without compromising accuracy.
-Figure 6: Module-specific Task-localization Results.
+Figure 6.
+Module-specific Task-localization Results.
 We plot the relationship between the accuracy and speed up of each module, where each diagram corresponds to the categorized module.
 The red-dotted line is the baseline performance (Prompt-tuning).
-5.5 Prior Knowledge Injection is Important
-Figure 7: Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
-This section describes the importance of injecting prior task knowledge (described Section 4.4) to accurately quantify the task relevance, as shown in Figure 7.
-We prune the top 50% of task neurons by using two methods (i.e., Skeleton and w/o Prior Task Knowledge) for a fair comparison.
-The experimental results reveal that prior task knowledge injection has a crucial role in quantifying the task relevance.
-Excluding the Prior Knowledge Injection step results in slower convergence and lower task accuracy compared to our method.
-5.6 Skeleton is Scalable to a Larger Model
-We conduct experiments to verify that our method is scalable to a larger model, as shown in Table 2.
-Specifically, we adopt BERT-large to show the scalability of our framework.
-These results reveal that our method is scalable to a larger model.
-Method SST-2 IMDB
-BERT-base (110M) Fine-tuning (Single) 92.39 92.50
-Fine-tuning (Multi) 91.36 92.16
-Prompt-tuning 89.36 88.56
-Skeleton 89.21 88.65
+5.5. Ablation Study on Knowledge Condensation
+We conduct the ablation study to justify the necessity of Knowledge Condensation, as described in The Knowledge Condensation Step (Section 4.4).
+Specifically, we eliminate neurons the same as the results of the DoE in the ablation study for a fair comparison, and do not conduct the knowledge condensation step.
+As a result, if we do not use Knowledge Condensation, the performance for SST-2 and IMDB significantly degrades.
+These results reveal the justification for adopting the knowledge condensation step.
+SST-2 IMDB
+DoE 89.21 88.65
+w/o Knowledge Condense 0.0 48.34
 […]
-[…]
-BERT-large (340M) Fine-tuning (Single) 92.77 93.57
-Fine-tuning (Multi) 92.73 93.83
-Prompt-tuning 91.28 89.96
-Skeleton 92.19 89.35
-[…]
-[…]
-Table 2: Experiments on various sizes of BERT.
-6 Conclusion
-In this study, we propose Skeleton, which explores the extreme of the prompt tuning method by localizing task-relevant neurons from the original language model.
-Skeleton investigates the optimal task-localized sub-network from a language model and condenses task knowledge by utilizing only task-relevant parameters during the prompt-tuning procedure.
-We demonstrate our method for various benchmarks and reveal that Skeleton dramatically enhances the efficiency of language models, showing significant inference speed boost-up.
-Limitations
-Our framework accelerates the inference speed of language models, identifying the optimal subnetwork for utilizing language models.
-However, while our approach exhibited comparable performance and occasionally slightly superior results to the existing prompt tuning method, these improvements were not consistent across overall task performance.
-In addition, our experiments are limited to five widely-used natural language understanding datasets; thus, we should extend to a wide range of datasets to validate the applicability of our framework.
-Furthermore, we select feed-forward networks to localize task neurons in language models since they are adequate for inference speed acceleration.
-However, the entire modules in Transformer should be considered as a localization target to achieve the extreme inference speed up; we leave this point as future research.
-Acknowledgements
-This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)].
+Table 4. Ablation Studies on Knowledge Condensation.
+5.6. Prior Knowledge Aligning is Important
+To accurately quantify task relevance, this section highlights the role of aligning prior task knowledge using prompt tuning, as described in The Knowledge Quantification Step (Section 4.4) and illustrated in Figure 7.
+We prune the top 50% of task neurons by using two methods (i.e., DoE and w/o Prior Task Knowledge) for a fair comparison.
+The experimental results reveal that prior task knowledge alignment has a crucial role in quantifying the task relevance.
+Excluding the Prior Knowledge Alignment results in slower convergence and lower task accuracy than our method.
+Figure 7. Prior Knowledge Injection is Important. We plot the mean accuracy (± one standard deviation for three trials) of BERT-base for SST-2 and IMDB.
+6. Conclusion
+In this study, we proposed Decomposition of Experts (DoE), a novel framework that optimizes pre-trained language models by identifying and activating task-specific experts.
+To the best of our knowledge, our method is the first to propose a novel usage paradigm for language models that temporarily deactivates irrelevant experts (i.e., neurons) and restores the full model to await subsequent tasks.
+DoE leverages an unplug-and-play strategy to enhance inference efficiency while maintaining task performance.
+Experimental results show that DoE achieves up to a 1.73x speed-up with a 65% parameter pruning rate across diverse tasks.
+Comparative evaluations and ablation studies validate the effectiveness of DoE and its components.
+The scalability of DoE makes it a practical and adaptable solution for transformer-based architectures.
+Acknowledgements.
+This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)].
 K.
 Jung is with ASRI, Seoul National University, Korea.
-Appendix A Instruction Templates.
-We use text templates for input texts of the utilized datasets.
-Specifically, the text templates consist of an instruction and an input text, as shown in Table 3.
-For the prediction, we utilize the representation of the mask token (i.e., "<mask>") to utilize the pre-trained knowledge of encoder-only language models.
-Appendix B Module-specific Task-localization Details
+The Institute of Engineering Research at Seoul National University provided research facilities for this work.
+GenAI Usage Disclosure
+Generative AI was not used at any stage of the research, except for minor corrections related to spelling or grammar in the writing.
+Appendix
+Appendix A Module-specific Task-localization Details
 We conduct experiments for various modules in the transformer architecture, as described in Section 5.4.
-We implement algorithms for pruning those networks following our Skeleton framework to satisfy the dimension of adjacent networks after pruning.
+We implement algorithms for pruning networks following our DoE framework to satisfy the dimension of adjacent networks after pruning.
 For example, suppose that we prune the Key network, $W^{K}\in\mathbb{R}^{d\times d}$, and get a pruned weight, $\tilde{W}^{K}\in\mathbb{R}^{d\times d^{\prime}}$, in the Attention module.
-Then, we should also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since they are adjacent networks.
-Therefore, after the pruning process, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by adding zero values to ensure the identical dimension.
-We also implement other layers in the same way with it.
+Then, we also prune the Query network, $W^{K}\in\mathbb{R}^{d\times d}$, since it is an adjacent network.
+Therefore, if we get $v_{Q}\in\mathbb{R}^{d}$ and $v_{K}\in\mathbb{R}^{d^{\prime}}$ from the Query and Key networks after pruning, we recover the pruned parameters of $v_{K}\in\mathbb{R}^{d^{\prime}}$ by inserting zero values to ensure the identical dimension.
+We also implement other layers in the same way.